OpenStudy (anonymous):

True or False!!! :) #1: sin(-x) = -cos x for all values of x. ***my answer; FALSE is that right? @terenzreignz :)

OpenStudy (anonymous):

#2: also true or false! :) sin(a+b) = sin a cos b - cos a sin b for all values of a and b. ***my answer: FALSE is that right? :)

terenzreignz (terenzreignz):

First off, nice new colour :D Now... let's see...

OpenStudy (anonymous):

color? :/ haha :/ lol confusion!

terenzreignz (terenzreignz):

Correct on both questions.

OpenStudy (anonymous):

yayyy!! :) thx!!! :D

terenzreignz (terenzreignz):

For future reference sin(-x) = -sin(x) NOT -cos(x) And sin(a+b) = sin(a)cos(b) + cos(a)sin(b) NOT sin(a)cos(b) - cos(a)sin(b)

OpenStudy (anonymous):

ahhh okay :) i see!! i'll keep that in mind!! Thanks!!! :D